# Operator Deep Smoothing for Implied Volatility

### v1

@@ 3.2 Operator Deep Smoothing
Let $\mathbf{v}=\{v(x_{l})\}_{l=1}^{p}$ be the collection of observed data, for example implied volatilities as in Section 2. This notation silently adopts the neural operator philosophy, connecting the data point $v(x_{l})$ with coordinates $x_{l}$, hinting at a latent function $v\colon D\to\mathbb{R}$ giving rise to the observed values. The smoothing or interpolation task consists of constructing an appropriately regular candidate $\hat{v}$ for $v$ from the data $\mathbf{v}=v|_{\pi}$, with $\pi=\{x_{1},\dots,x_{p}\}$. The operator deep smoothing approach uses a neural operator $\tilde{F}^{\theta}$, trained using historical data, to generate $\hat{v}:=\tilde{F}^{\theta}(\mathbf{v})$. It fundamentally leverages the discretization-invariance to produce consistent results even if the "sensors" $x_{l}$ of the latent function $v$ continuously change in availability and/or location in the domain $D$. This is the situation for volatility smoothing, where the sensors $x_{l}=(\tau_{l},k_{l})$ are the (time-to-expiry, log-moneyness) coordinates of the quoted options, which – as noted in Section 1 and illustrated in Figure 1(b) – move with the market, thus motivating the operator deep smoothing approach. We now describe the methodology in more detail. For a first glance at the results we refer to Figure 2.
@@
Depending on the application, the smoothing task may be subject to constraints. For volatility smoothing, the smoothed surface $\hat{v}^{\theta}=\tilde{F}^{\theta}(\mathbf{v})$ must be free of static arbitrage. This is effectively enforced by augmenting the loss function with additional penalization terms, moving away from a pure operator learning problem. [^5] It is not the goal of volatility smoothing to learn the identity operator throughout the entire input space. This does not only promote the relevant properties in the neural operator output but can also help define it when faced with sparsity of data in the domain $D$ (in this context see also [23]). From Theorem 2.1, these penalization terms naturally motivate our choices of $\mathcal{L}_{\text{but}}$ and $\mathcal{L}_{\text{cal}}$. The strike arbitrage constraint 2.1(ii) is handled via
@@
where we ignore the asymptotic condition since our experiments are focused on the bounded domain $D$ (Figure 1(a)). The inclusion of $\varepsilon$ promotes strictly positive implied densities (we will use $\varepsilon=10^{-3}$), while we choose the $1$-norm to induce sparsity in the constraint violation. The calendar arbitrage constraint 2.1(i) can be tackled analogously with
@@
Interpolating Neural Operator Architectures
Various neural operator architectures exist, mostly arising from the kernel integral transform framework of [20]. Most prominently, these include Fourier neural operators (FNO), delivering state-of-the-art results on fixed grid data, as well as graph neural operators (GNO), able to handle arbitrary mesh geometries, both reviewed in Appendix A.1. While highly effective with documented universality, these neural operators are not directly applicable for interpolation as their layers include a pointwise-applied linear transformation, which limits the output to the set of the input data locations. Dropping this local linear transformation results in an architecture proved to retain universality [20] and – at least for its implementation as a GNO – fit for interpolation tasks. In fact, universality holds [21] for an architecture combining the local linear transformation with a simple averaging operation, suggesting the fundamental importance of the collaboration of local and non-local components; this was noted in [20] for whom retaining the local components can be "beneficial in practice". We confirm this in our experiments showing that a purely non-local architecture has substantially reduced performance.
We therefore propose a new architecture for operator deep smoothing leveraging GNOs’ unique ability to handle irregular mesh geometries. We use a purely non-local first layer (dropping the pointwise linear transformation), and use it to produce hidden states at all required output locations, enabling subsequent layers to retain their local transformations. Since GNOs do not theoretically guarantee a smooth output, we augment the training with additional regularization terms such as $\mathcal{L}_{\text{reg}}(\theta;\mathbf{v})=\|\Delta\hat{v}^{\theta}\|_{2}$, with $\Delta$ the Laplace operator, and provide a full description in Appendix B.
@@ 4.1 Model Training
Motivated by Figure 1(a), we transform the time-to-expiry and log-moneyness coordinates via $\rho=\sqrt{\tau}$ and $z=k/\rho$. Intuitively, this transformation converts the "natural" scaling of implied volatility by time-to-expiry to a scaling of the input domain. From here on, we consider the domain in these coordinates, setting $D=(\rho_{\text{min}},\rho_{\text{max}})\times(z_{\text{min}},z_{\text{max}})=(0.01,1)\times(-1.5,0.5)$. In $(\tau,k)$-coordinates, $D$ becomes a cone-shaped region, that, on average, contains 96.6% of traded options (with time-to-expiry below one year) and, with respect to $(\rho,z)$-coordinates, is more evenly populated, improving the numerics.
@@
The model hyperparameters (giving rise to 102529 trainable parameters in total), identified by manual experimentation, are detailed in Appendix C.2.
@@
We implement a Vega-weighted (see (2)) version of the fitting loss $\mathcal{L}_{\text{fit}}$ (9). We compute $\mathcal{L}_{\text{but}}$ directly as (10) (in the transformed coordinates) on a synthetic grid using finite differences. For $\mathcal{L}_{\text{cal}}$, we implement a multiplicative version that is invariant to the scale of implied volatility. We provide a precise description in Appendix C.3.
@@
We train the GNO $F^{\theta}$ over $500$ epochs on the training dataset using the AdamW optimizer with learning rate $\lambda=10^{-4}$ and weight decay rate $\beta=10^{-5}$, and use a pseudo batch size of 64 by accumulating gradients. We randomly sub-sample the inputs $\mathbf{v}$ and randomize the grids on which we compute the arbitrage losses. The training is performed in around 200 hours using a NVIDIA Quadro RTX 6000 GPU. The validation loss is reported in Appendix C.3 in Table 1.
@@ 4.2 Results
Figure 3: Benchmark metrics (surface-averages, between 25%- and 75%-quantiles) over training period (computed from $\mathcal{D}_{\text{val}}$) and testing period, as well as surface-averaged spread. Resampled, monthly for training and weekly for testing period.
@@
Model Finetuning
During production use, the GNO would be retrained regularly using the most recent available data. We emulate this procedure by evaluating the benchmark metrics on a monthly basis over the 2021 test dataset $D_{\text{test}}$: Following the evaluation of the first month’s test data, the GNO is trained for 10 epochs on the test data just reviewed, with each mini-batch augmented by an equal amount of data from the training dataset $\mathcal{D}_{\text{train}}$. We repeat this, progressively incorporating an additional month of data, until the entire dataset $\mathcal{D}_{\text{test}}$ is assessed. This finetuning-evaluation procedure takes circa 1.8 GPU hours per month.
@@
It is apparent from Figure 3 that the operator deep smoothing approach substantially improves on SVI’s smoothing capabilities, with respect to both $\delta_{\text{abs}}$ and $\delta_{\text{spr}}$. [^7] We produce the SVI benchmark as described in Section E. The use of our approach in practice, with monthly finetuning, can be expected to smooth the volatility surface with a MAPE of around 0.5%, while SVI fluctuates between 1% and 2%. Our approach is competitive with [2], which perform instance-by-instance volatility smoothing using neural networks, and report a MAPE of around 1% for the period Jan-Apr 2018.888However, [2] does not perform a similar restriction of the domain of the volatility surface. The various figures in Appendix D show the qualitative improvements of our method versus the SVI benchmark.
An important metric to consider when making sense of Figure 3 is the historical tightening of the bid-ask spread, driven by increases in competition on the S&P 500 option market over recent years. This fact explains why $\delta_{\text{spr}}$ is very small on the chronologically earlier training dataset, both for operator deep smoothing and SVI, even as $\delta_{\text{abs}}$ is comparatively larger. The latter fact is explained similarly (and is visually recognizable by Figure 8): Wide spreads give rise to more variable prices, leading to greater need for correction by the smoothing algorithm, captured by $\delta_{\text{abs}}$.
Complementary to Figure 3, Figure 4 resolves the error metrics as well as the terms controlling the absence of arbitrage spatially, averaged along the time axis. The MAPE tends to be larger on the Call wing, in accordance with noisier option prices (Call options experience less trading than Put options). Moreover, we discern that, on average, the smoothed surfaces are completely free of arbitrage.
@@ 5 Discussion
We provide a novel method for implied volatility smoothing, resulting from an application of our general operator deep smoothing approach for discretization-invariant data interpolation. The approach leverages a graph neural operator to directly map given data – consistently across data dimensions and spatial arrangement – to smoothed surfaces, transcending classical parametric smoothing techniques. In the example of volatility smoothing, benefits include a massively simplified online calibration process.
@@
Due to the discretization-invariance of the GNO architecture underlying the operator deep smoothing approach, the smoothing method is robust with respect to subsampling of inputs (we include this technique into operator training to improve generalization, compare Appendix C.3). In practice, subsampling of inputs occurs in the context of outlier removal. In the example of volatility smoothing, certain quotes may be determined spurious. Simply removing anomalous datapoints from the input is compatible with our method.
@@
Figure 3 makes the compression qualities of the operator deep smoothing approach apparent: We compute the entire historical timeseries using a single GNO instance, with around 100 thousand parameters. Evaluating the SVI benchmark, on the other hand, requires 61454 model instances (one per slice), or a total of 307270 parameters. A comparison with [2], which for each smoothed surface instantiates a deep neural network corrector of around 5085 parameters, [^9] Computed as the sum of $120=3\times 40$ parameters for the input layer, three times $1640=41\times 40$ parameters for the hidden layers, $41$ parameters for the output layers, plus $4$ additional parameters of the SSVI prior and a scaling parameter. is striking. Performing [2]’s suggested smoothing on the entire dataset 2012–2021 would require more than 200 million parameters. On the other hand, we expect our operator deep smoothing approach to deliver competitive results continuously over the entire training period and beyond (with regular finetuning), even when moving to higher-frequency data (our data is summarized on a 20-minute interval basis).
@@
Compared to ad-hoc volatility parametrizations like SVI, the operator deep smoothing approach loses interpretability of parameters. Depending on the application, the computation of sensitivities with respect to intuitive parameters may be a stringent requirement. This disadvantage is generally shared by neural network based models in engineering applications.
More generally, in some situations dimensionality reduction (even without interpretability of parameters) may be a desirable additional feature that is not directly achieved by our operator deep smoothing approach. Combining the VAE method [5] with our operator deep smoothing approach could lead to further promising potential applications of neural operators. [18] introduces neural mappings, which generalize neural operators to mixed infinite-/finite-dimensionality for input or output spaces. This motivates a discretization-invariant GNO-based encoder, fit to handle raw incoming market data, and a classical decoder to extend the operator deep smoothing approach to a VAE-like architecture.
@@ Acknowledgements
Access to data was made possible through the Clearify Project funding by an Imperial College SME engagement grant jointly with Zeliade Systems, as well as the MSc in Mathematics and Finance, Imperial College London. AJ gratefully acknowledges financial support from the EPSRC grants EP/W032643/1 and EP/T032146/1. Computational resources and support were provided by the Imperial College Research Computing Service [14]. For the purpose of open access, the author(s) has applied a Creative Commons Attribution (CC BY) licence (where permitted by UKRI, ‘Open Government Licence’ or ‘Creative Commons Attribution No-derivatives (CC BY-ND) licence’ may be stated instead) to any Author Accepted Manuscript version arising.

### v2

@@ 3.2 Operator Deep Smoothing
Let $\mathbf{v}=\{v(x_{l})\}_{l=1}^{p}$ be the collection of observed data, for example implied volatilities as in Section 2. This notation silently adopts the neural operator philosophy, connecting the data point $v(x_{l})$ with coordinates $x_{l}$, hinting at a latent function $v\colon D\to\mathbb{R}$ giving rise to the observed values. The smoothing or interpolation task consists of constructing an appropriately regular candidate $\hat{v}$ for $v$ from the data $\mathbf{v}=v|_{\pi}$, with $\pi=\{x_{1},\dots,x_{p}\}$. The operator deep smoothing approach uses a neural operator $\tilde{F}^{\theta}$, trained using historical data, to generate $\hat{v}:=\tilde{F}^{\theta}(\mathbf{v})$. It fundamentally leverages the discretization-invariance to produce consistent results even if the ”sensors” $x_{l}$ of the latent function $v$ continuously change in availability and/or location in the domain $D$. This is the situation for volatility smoothing, where the sensors $x_{l}=(\tau_{l},k_{l})$ are the (time-to-expiry, log-moneyness) coordinates of the quoted options, which – as noted in Section 1 and illustrated in Figure 1(b) – move with the market, thus motivating the operator deep smoothing approach. We now describe the methodology in more detail. For a first glance at the results we refer to Figure 2.
@@
Depending on the application, the smoothing task may be subject to constraints. For volatility smoothing, the smoothed surface $\hat{v}^{\theta}=\tilde{F}^{\theta}(\mathbf{v})$ must be free of static arbitrage. This is effectively enforced by augmenting the loss function with additional penalization terms, moving away from a pure operator learning problem. [^5] It is not the goal of volatility smoothing to learn the identity operator throughout the entire input space. This does not only promote the relevant properties in the neural operator output but can also help define it when faced with sparsity of data in the domain $D$ (in this context see also Li et al., (2021)). From Theorem 2.1, these penalization terms naturally motivate our choices of $\mathcal{L}_{\text{but}}$ and $\mathcal{L}_{\text{cal}}$. The strike arbitrage constraint 2.1(i) is handled via
@@
where we ignore the asymptotic condition since our experiments are focused on the bounded domain $D$ (Figure 1(a)). The inclusion of $\varepsilon$ promotes strictly positive implied densities (we will use $\varepsilon=10^{-3}$), while we choose the $1$-norm to induce sparsity in the constraint violation. The calendar arbitrage constraint 2.1(ii) can be tackled analogously with
@@
Interpolating Graph Neural Operator
Various neural operator architectures exist, mostly arising from the kernel integral transform framework of Kovachki et al., (2023). Most prominently, these include Fourier neural operators (FNO), delivering state-of-the-art results on fixed grid data, as well as graph neural operators (GNO), able to handle arbitrary mesh geometries (both reviewed in Appendix A.1). While highly effective with documented universality, these neural operators are not directly applicable for interpolation tasks as their layers include a pointwise-applied linear transformation, which limits the output to the set of the input data locations. Dropping this local transformation results in an architecture proved to retain universality Kovachki et al., (2023) and that – at least for its implementation as a GNO – allows to interpolate functions. On the other hand, Lanthaler et al., (2023) proves universality for the architecture combining the local linear transformation with a simple averaging operation, suggesting the fundamental importance of the collaboration of local and non-local components for the expressivity of neural operators. This was noted in Kovachki et al., (2023), for whom retaining the local components can be “beneficial in practice”, and confirmed in our experiments, where a purely non-local architecture led to substantially reduced performance.
We therefore propose a new architecture for operator deep smoothing leveraging GNOs’ unique ability to handle irregular mesh geometries. We use a purely non-local first layer (dropping the pointwise linear transformation), and use it to produce hidden state at all required output locations, enabling subsequent layers to retain their local transformations. Since GNOs do not theoretically guarantee a smooth output, we augment the training with additional regularization terms such as $\mathcal{L}_{\text{reg}}(\theta;\mathbf{v})=\|\Delta\hat{v}^{\theta}\|_{2}$, with $\Delta$ the Laplace operator, and provide a full description in Appendix B.
@@ 4.1 Model Training
Motivated by Figure 1(a), we transform time-to-expiry and log-moneyness via $\rho=\sqrt{\tau}$ and $z=k/\rho$. Intuitively, this transformation converts the“natural” scaling of implied volatility by the square root of time-to-expiry to a scaling of the input domain. From here on, we consider the domain in these coordinates, setting $D=(\rho_{\text{min}},\rho_{\text{max}})\times(z_{\text{min}},z_{\text{max}})=(0.01,1)\times(-1.5,0.5)$. In $(\tau,k)$-coordinates, $D$ becomes a cone-shaped region, that, on average, contains 96.6% of traded options (with time-to-expiry below one year) and, with respect to $(\rho,z)$-coordinates, is more evenly populated, improving the numerics.
@@
We remind that we rely on the interpolating graph neural operator introduced in Section 3.2 and described in detail in Appendix B. The model hyperparameters (giving rise to 102529 trainable parameters in total) were identified by manual experimentation and are detailed in Appendix C.2.
@@
We implement a Vega-weighted (see Appendix C.3) version of the fitting loss $\mathcal{L}_{\text{fit}}$ from (8). We compute $\mathcal{L}_{\text{but}}$ directly as (9) (in the transformed coordinates) on a synthetic grid using finite differences. For $\mathcal{L}_{\text{cal}}$, we implement a multiplicative version that is invariant to the level of implied volatility. We provide a precise description in Appendix C.3.
@@
We train the GNO over $500$ epochs on the training dataset using the AdamW optimizer with learning rate $\lambda=10^{-4}$ and weight decay rate $\beta=10^{-5}$, and use a pseudo batch size of 64 by accumulating gradients. We randomly sub-sample the inputs $\mathbf{v}$ and randomize the grids on which we compute the arbitrage losses. The training is performed in around 200 hours using a NVIDIA Quadro RTX 6000 GPU. The validation loss is reported in Appendix C.3 in Table 2. Code and trained model weights are provided in the supplementary material.
@@ 4.2 Results
Figure 3: Benchmark metrics (surface-averages, between 25%- and 75%-quantiles) over training period (computed from $\mathcal{D}_{\text{val}}$) and testing period, as well as surface-averaged spread. Resampled, monthly for training period and weekly for testing period.
@@
Evaluation and Model Finetuning
During production use, the GNO would be retrained regularly using the most recent available data. We emulate this procedure during our evaluation of the benchmark metrics over the test dataset $\mathcal{D}_{\text{test}}$ (containing the year 2021): Following the evaluation of the first month’s test data, the GNO is trained for 10 epochs on the test data just reviewed, with each mini-batch augmented by an equal amount of data from the training dataset $\mathcal{D}_{\text{train}}$. We repeat this, progressively incorporating an additional month of data, until the entire dataset $\mathcal{D}_{\text{test}}$ is assessed. This finetuning-evaluation procedure takes circa 1.8 GPU hours per month.
@@
It is apparent from Figure 3 that our operator deep smoothing approach substantially improves on SVI’s smoothing capabilities, with respect to both $\delta_{\text{abs}}$ and $\delta_{\text{spr}}$. [^7] We produce the SVI benchmark as described in Section E. Our approach, with monthly finetuning, smooths the volatility surface with a MAPE of around 0.5%, while SVI fluctuates between 1% and 2%. The various figures in Appendix D illustrate the qualitative improvements of our method versus SVI. Moreover, our approach appears highly competitive with Ackerer et al., (2020), which performs instance-by-instance volatility smoothing using classical neural networks and reports a MAPE of around 1% for synthetically generated data. [^8] Compare Table 1 of Ackerer et al., (2020). We note that Ackerer et al., (2020) does not perform a similar restriction of the domain of the volatility surface. We reenact its backtesting for the period Jan-Apr 2018 using our method and summarize the results in Table 5 in Appendix C.5.999We emphasize the following aspects. First, the backtest involves dropping half of input datapoints for each surface, and our trained operator continues to perform accurate smoothing, a strong indicator of the robustness of our approach with respect to subsampling of inputs afforded by neural operator’s discretization invariance. Second, while Ackerer et al., (2020) requires to train 61 neural networks to perform the backtest once, our operator approach enables us to run 25 repetitions in around two minutes on a consumer grade laptop CPU, which is the average time that it takes Ackerer et al., (2020) to train one network.
An important fact to consider when analyzing Figure 3 is the historical tightening of the bid-ask spread (displayed in the third row), driven by increases in competition on the S&P 500 option market. This explains why $\delta_{\text{spr}}$ is very small early in the training dataset, both for operator deep smoothing and SVI, while $\delta_{\text{abs}}$ is large: Wide spreads make $\delta_{\text{spr}}$ more lenient an error metric but are accompanied by noisier prices, necessitating greater need for correction by the smoothing algorithm, in turn captured by $\delta_{\text{abs}}$. [^10] This is visually discernible from a comparison of Figure 8 and Figure 10. This argumentation extends to spikes in the spread, which indicate periods of market stress, and then allows to explain spikes in $\delta_{\text{abs}}$.
Complementary to Figure 3, Figure 4 resolves the error metrics as well as the terms controlling the absence of arbitrage spatially, averaged over time. $\delta_{\text{abs}}$ tends to be larger on the Call side (positive log-moneyness), in accordance with Call option’s noisier prices (Call options experience less trading than Put options). Moreover, we discern that, on average, the smoothed surfaces are completely free of arbitrage (indicated by non-negativity).
@@
Generalization
To test the generalization capabilities of our approach, we procure end-of-day options data for the S&P 500 (SPX), the NASDAQ-100 (NDX), the Dow Jones Industrial Average (DJX), and the Russell 2000 (RUT) for the month of January 2021 from the OptionMetrics Ivy DB US database, accessed by us through the Wharton Research Data Services (WRDS). We evaluate the trained operator on the data for the month of January 2021 (right after the training period on the S&P 500 intraday data), and report the average error metrics $\delta_{\text{abs}}$ and $\delta_{\text{spr}}$ as well as the average arbitrage losses $\mathcal{L}_{\text{cal}}$ and $\mathcal{L}_{\text{but}}$ in Table 1. Firstly, our method maintains its performance on end-of-day S&P 500 data, validating the soundness of our approach: While end-of-day data is slightly different from intraday data, our method still yields small error metrics and arbitrage-free prices. Secondly, the method generalizes well to other indices. We want to stress the fact that our operator has solely been trained on intraday S&P 500 data. Its accurate and virtually arbitrage-free output on end-of-day data of other indices is a strong indicator of the robustness of our approach. We provide further example plots for these datasets in Appendix D.3.
Table 1: Average error metrics and arbitrage losses for end-of-day options data for US indices in January 2021. The GNO has been trained solely on intraday S&P 500 data from before 2021.
SPX NDX DJX RUT
$\delta_{\text{abs}}$ 0.00272 0.01057 0.01629 0.00885
$\delta_{\text{spr}}$ 0.64423 1.53306 0.20736 1.03183
$\mathcal{L}_{\text{but}}$ 6.41e-05 2.16e-04 1.61e-03 4.48e-06
$\mathcal{L}_{\text{cal}}$ 0.00000 0.00000 3.13e-08 0.00000
@@ 5 Discussion
We provide a novel method for implied volatility smoothing, resulting from an application of our general operator deep smoothing approach for discretization-invariant data interpolation. The approach leverages a graph neural operator to directly map given data – consistently across size and spatial arrangement – to smoothed surfaces, transcending classical parametric smoothing techniques. In the example of volatility smoothing, benefits include a massively simplified online calibration process.
Learning from Large Datasets
By moving the application of neural networks from the instance-by-instance level (Ackerer et al.,, 2020) to the “operator level”, we leverage the information contained in the entire training dataset for the smoothing of every single surface. In other words, our method “unlocks” large historical options datasets for volatility smoothing. We argue that our substantial outperformance against Ackerer et al., (2020) in the “Extrapolation-Test”-setting of the benchmark detailed in Table 5 of Appendix C.5 owes to this circumstance.
@@
The discretization-invariance of the GNO entails that our method is robust with respect to subsampling of inputs. In practice, subsampling of inputs occurs in the context of outlier removal. In the example of volatility smoothing, certain quotes may be determined spurious. Simply removing anomalous datapoints from the input is compatible with our method (moreover, we leverage this fact during operator training to improve generalization, compare Appendix C.3).
@@
Figure 3 makes the compression qualities of the operator deep smoothing approach apparent: We compute the entire historical timeseries using a single GNO instance, with around 100 thousand parameters. Evaluating the SVI benchmark, on the other hand, requires 61454 model instances (one per slice), or a total of 307270 parameters. A comparison with Ackerer et al., (2020), which for each smoothed surface trains a new neural network of around 5085 parameters, [^11] Computed as the sum of $120=3\times 40$ parameters for the input layer, three times $1640=41\times 40$ parameters for the hidden layers, $41$ parameters for the output layers, plus $4$ additional parameters of the SSVI prior and a scaling parameter. is striking: Smoothing of the CBOE dataset 2012–2021 at its 20-minute interval frequency would require more than 200 million parameters (more with rising frequency). At the same time, we expect our GNO to perform accurate smoothing over the entire training period and beyond (with regular finetuning), and our model instance remains fixed, even when moving to higher-frequency data.
@@
Compared to ad-hoc volatility parametrizations like SVI, the operator deep smoothing approach loses interpretability of parameters, which for some practitioners may be a stringent requirement. This disadvantage is generally shared by neural network based engineering solutions. Moreover, in some situations dimensionality reduction (even without interpretability of parameters) may be a desirable additional feature that is not directly achieved by our operator deep smoothing approach. Combining the VAE method (Bergeron et al.,, 2021) with our operator deep smoothing approach could lead to further promising potential applications of neural operators. Huang et al., (2024) introduces neural mappings, which generalize neural operators to mixed infinite-/finite-dimensionality for input or output spaces. This motivates a discretization-invariant GNO-based encoder, fit to handle raw incoming market data, and a classical decoder to extend the operator deep smoothing approach to a VAE-like architecture.
@@ Acknowledgements
Access to data was made possible through the Clearify Project funding by an Imperial College SME engagement grant jointly with Zeliade Systems, as well as the MSc in Mathematics and Finance, Imperial College London. AJ gratefully acknowledges financial support from the EPSRC grants EP/W032643/1 and EP/T032146/1. Computational resources and support were provided by the Imperial College Research Computing Service (Harvey,, 2017). For the purpose of open access, the author(s) has applied a Creative Commons Attribution (CC BY) licence (where permitted by UKRI, ‘Open Government Licence’ or ‘Creative Commons Attribution No-derivatives (CC BY-ND) licence’ may be stated instead) to any Author Accepted Manuscript version arising.